# One-Layer Transformers are Provably Optimal for In-context Reasoning and Distributional Association Learning in Next-Token Prediction Tasks

### v2

@@ Appendix B Missing Proofs in Section 4
B.2 Analysis of Normalized Gradient Descent on Population and Empirical Losses
@@
B.2.1 Known $\alpha$: Running Normalized Gradient Descent on the Population Loss $L(\lambda)$
@@
By Lemma C.1, the derivative of $f(\lambda)=-C\lambda+\ln(\frac{e^{C\lambda}}{1-\alpha}+N-1)$ is negative. Hence, $\frac{dL_{\mathrm{pop}}}{d\lambda}<0$. It follows that running normalized gradient descent on $L_{\mathrm{pop}}(\lambda)$ from $\lambda_{0}=0$ gives
@@
B.2.2 Unknown $\alpha$: Proof Of Theorem 4.3
@@
By Lemma C.1, we have $\frac{dL_{\mathrm{emp}}}{d\lambda}<0$. As a result, running normalized gradient descent on $L_{\mathrm{emp}}$ gives $\lambda_{t}=\eta t$. The population loss is
@@
Appendix C Technical Lemmas
Lemma C.1.
@@ Appendix C Technical Lemmas
Appendix D Further Details on Experiments
D.1 Additional Details on the Experimental Setup
@@
We also experimented with several different values of learning rates ranging from $0.05$ to $0.5$. Consistent with the theoretical findings, we find that the more reparamterized a model is, the less sensitive it is to changes in the learning rate. All of our results are reported for learning rates set at either $0.1$ or $0.5$.
@@
Computing Resources The experiments are implemented in PyTorch. All experiments are run on a single-CPU computer. The processor is 11th Gen Intel(R) Core i7-11700K with $32$ GB RAM. Training each model takes about $30$ minutes from start to finish.
D.2 Additional Results
D.2.1 Minimizing Population Loss
Figure 3 to Figure 9 show results for models that were not reported in Figure 1. For each model, we show how population loss and unseen output test loss converge as a function of the number of iterations. For some models such as Origin-Softmax-FA, larger learning rates are better and seem to allow the model to escape bad saddle points. On the contrary, for other models such as Origin-ReLU-F, larger learning rates lead to more unstable training dynamics. For all fully-reparameterized models, larger learning rates lead to faster convergence, which is consistent with our theoretical results.
Figure 3: Population and Unseen Output Test Losses of Origin-Softmax-FA with $\eta=0.1$ (first row), $\eta=0.1$ (second row). The population losses do not converge at $\eta=0.1$ but converges at $\eta=0.5$. Generalization to unseen output words is non-existent.
Figure 4: Population and Unseen Output Test Losses of Origin-Linear-F with $\eta=0.1$ (first row), $\eta=0.1$ (second row). The population losses do not converge at neither $\eta=0.1$ nor $\eta=0.5$. Generalization to unseen output words is non-existent.
Figure 5: Population and Unseen Output Test Losses of Origin-ReLU-F with $\eta=0.1$ (first row), $\eta=0.5$ (second row). The population losses converge at $\eta=0.1$ but do not converge at $\eta=0.5$. Generalization to unseen output words is non-existent.
Figure 6: Population and Unseen Output Test Losses of Reparam-Softmax-F with $\eta=0.1$ (first row), $\eta=0.5$ (second row). The population losses do not converge at neither $\eta=0.1$ nor $\eta=0.5$. The model has a high degree of generalization to unseen output words.
Figure 7: Population and Unseen Output Test Losses of Reparam-ReLU-F with $\eta=0.1$ (first row), $\eta=0.5$ (second row). The population loss for noiseless setting converges at both $\eta=0.1$ nor $\eta=0.5$. All noisy population losses never converge for any learning rate. The model has a high degree of generalization to unseen output words.
Figure 8: Population and Unseen Output Test Losses of Reparam-Linear-F with $\eta=0.1$ (first row), $\eta=0.5$ (second row). The population loss for noiseless setting converges at both $\eta=0.1$ nor $\eta=0.5$. All noisy population losses never converge for any learning rate. The model has a high degree of generalization to unseen output words.
@@ D.2.1 Minimizing Population Loss
Figure 9: Population and Unseen Output Test Losses of Reparam-Linear-W-FA with $\eta=0.1$ (first row), $\eta=0.5$ (second row). Population losses converge with limited generalization to unseen output words.
@@
D.2.2 Attention Layer Learns to Predict Output Tokens while Feed-Forward Layer Learns to Predict Noise Token
@@
To measure the extent to which we can separate the learning functionality of the attention layer and the feed-forward layer, we train three models Origin-Linear-FA, Reparam-Linear-FA and Reparam-Linear-W-FA, and record the logits of each layer on the output tokens, the noise tokens and the maximum values in the logits of the two layers on all three noisy tasks with $\alpha=0.2,0.5$ and $0.8$. We use $\eta=0.1$ in all experiments. The results are reported in Figures 10 to 12.
@@
We say that the attention layer and the feed-forward layer learns to predict the output and noise tokens, respectively, if $\xi_{A,y}=\max_{j}\xi_{A,j}$ and $\xi_{F,\tau}=\max_{j}\xi_{F,\tau}$. It can be observed that all three models exhibit some layer-specific learning mechanism, including the original model where all three matrices ${\bm{V}},{\bm{W}}$ and $\bm{F}$} are trained from scratch without reparameterization. However, depending on the noise level, at least one of the two layers in the original model do not fully specialize in either the output nor the noise tokens. For $\alpha\leq 0.5$, Figures 10(a) and 10(d) show that the attention layer in Origin-Softmax-FA learns to predict output tokens perfectly, however the feed-forward layer does not always predict $\tau$. At $\alpha=0.8$, the feed-forward layer succeeds in learning to predict $\tau$ but the attention layer fails to focus entirely on the output tokens.
@@
In contrast, the fully-reparameterized model Reparam-Linear-FA exhibits perfect separation in the functionality of the two layers, which verified our Theorem 4.5. The same phenomenon is also observed in Reparam-Linear-W-FA, which suggests that reparameterizing $\bm{F}$} is sufficient to force the two layers to be biased towards two different types of tokens.
@@
Figure 10: Origin-Linear-FA with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.
@@
Figure 11: Reparam-Linear-W-FA with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.
@@
Figure 12: Reparam-Linear-FA with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.

### v3

@@ Appendix D Missing Proofs in Section 5
Lemma D.2.
(Optimality of softmax attention for noisy task) By setting ${\bm{U}}=[E(1)~E(2)~\dots~E(N)~E(N+1)]^{\top},{\bm{V}}=s{\bm{I}}_{d},{\bm{W}}=\lambda\sum_{q\in{\mathcal{Q}}}E(q)(\tilde{E}^{\top}(q)-2E^{\top}(\tau)-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top})$ and ${\bm{F}}=E(\tau)\sum_{q\in{\mathcal{Q}}}(\gamma E^{\top}(q)+\tilde{E}^{\top}(q))$, for all $\alpha\in(0,1)$, using softmax attention we obtain
$\lim_{s\to\infty,\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}L(s,\gamma,\lambda)\vcentcolon$= (76)
$\lim_{s\to\infty}\lim_{\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}\mathbb{E}_{y,z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\right]=L_{\mathrm{Bayes}}.$ (77)
Proof.
Consider a sentence with a trigger $q$ and output $y$. We have
${\bm{x}}_{H}^{\top}{\bm{W}}$ $=(E(q)+\tilde{E}(z_{H-1}))^{\top}\lambda\sum_{q^{\prime}\in{\mathcal{Q}}}E(q^{\prime})(\tilde{E}^{\top}(q^{\prime})-2E^{\top}(\tau)-\sum_{x=1,x\neq q^{\prime}}^{N}\tilde{E}(x)^{\top})$ (78)
$=\lambda(\tilde{E}^{\top}(q)-2E^{\top}(\tau)-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top}).$ (79)
It follows that
${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ $=(E(q)+\tilde{E}(z_{H-1}))^{\top}\lambda E(q)\left(\tilde{E}^{\top}(q)-2E^{\top}(\tau)-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top}\right){\bm{x}}_{h}$ (80)
$=\lambda\left(\tilde{E}^{\top}(q)-2E^{\top}(\tau)-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top}\right)(E(z_{h})+\tilde{E}(z_{h-1}))$ (81)
$=\lambda\left(-2\mathbb{1}\{z_{h}=\tau\}+\mathbb{1}\{z_{h-1}=q\}-\mathbb{1}\{z_{h-1}\neq q\}\right).$ (82)
Hence,
${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}=\begin{cases}\lambda&\textit{if }z_{h-1}=q,z_{h}=y\\ -\lambda&\textit{if }z_{h-1}=q,z_{h}=\tau\\ -\lambda&\text{otherwise}.\end{cases}$ (83)
Consequently,
$\sum_{j=1}^{H}\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ $=\sum_{z_{j-1}=q,z_{j}=y}\exp(\lambda)+\sum_{(z_{j-1},z_{j})\neq(q,y)}\exp(-\lambda)$ (84)
$=C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)$ (85)
The attention score at the $h$-th token in a sentence is
$\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ $=\frac{\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})}{\sum_{j=1}^{H}\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{j})}$ (86)
$=\begin{cases}&\frac{\exp(\lambda)}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda)}\quad\text{if }z_{h-1}=q,z_{h}=y\\ &\frac{\exp(-\lambda)}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda)}\quad\text{otherwise.}\end{cases}$ (87)
Next, we compute $\xi_{A,j}$. With $j=y$, we have
$\xi_{A,y}$ $={\bm{e}}_{y}^{\top}{\bm{U}}{\bm{V}}\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}=s\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\}$ (88)
$=s\left(\sum_{z_{h-1}=q}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\}+\sum_{z_{h-1}\neq q}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=y\}\right)$ (89)
$=s\frac{C_{q,y}\exp(\lambda)+(C_{y}-C_{q,y})\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}.$ (90)
With $j\neq y$, we have
$\xi_{A,j}$ $={\bm{e}}_{j}^{\top}{\bm{U}}{\bm{V}}\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}=s\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=j\}$ (91)
$=s\sum_{z_{h-1}\neq q}\frac{\exp(-\lambda)}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda)}\mathbb{1}\{z_{h}=j\}$ (92)
$=s\frac{C_{j}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda_{q})}.$ (93)
Next, the logits of the feed-forward layer is
$\xi_{F}={\bm{U}}{\bm{F}}\left({\bm{x}}_{H}+\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{V}}{\bm{x}}_{h}\right)$
$={\bm{U}}E(\tau)\left(\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma E^{\top}(q^{\prime})+\tilde{E}^{\top}(q^{\prime})\right)(E(q)+\tilde{E}(z_{H-1})+s\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h})$
$={\bm{e}}_{\tau}\left(\gamma+s\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})E^{\top}(q^{\prime}){\bm{x}}_{h}+\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\tilde{E}^{\top}(q^{\prime}){\bm{x}}_{h}\right)$
$={\bm{e}}_{\tau}\left(\gamma+s\sum_{q^{\prime}\in{\mathcal{Q}}}\gamma\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h}=q^{\prime}\}+\sum_{h=1}^{H}\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})\mathbb{1}\{z_{h-1}=q^{\prime}\}\right)$
$={\bm{e}}_{\tau}\left(\gamma+\left(s\sum_{q^{\prime}\in{\mathcal{Q}}}\frac{\gamma C_{q^{\prime}}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}\right)\right)$
$\quad+{\bm{e}}_{\tau}\left(s\frac{C_{q,y}\exp(\lambda)+C_{\tau}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}+\sum_{q^{\prime}\in{\mathcal{Q}},q^{\prime}\neq q}\frac{C_{q^{\prime}}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}\right),$
where we used $E^{\top}(q^{\prime}){\bm{x}}_{h}=\mathbb{1}\{z_{h}=q^{\prime}\}$ and $\tilde{E}^{\top}(q^{\prime}){\bm{x}}_{h}=\mathbb{1}\{z_{h-1}=q^{\prime}\}$. As a result, the combined logits of attention and feed-forward layers is
  • If $j=y$, then
    $\xi_{y}$ $=\xi_{A,y}+\xi_{F,y}=\xi_{A,y}$ (94)
    $=s\frac{C_{q,y}\exp(\lambda)+(C_{y}-C_{q,y})\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}.$ (95)
    It follows that $\lim_{\lambda\to\infty}\xi_{y}=s$.
  • If $j=\tau$, then
    $\xi_{\tau}$ $=\xi_{A,\tau}+\xi_{F,\tau}$ (96)
    $=s\frac{2C_{\tau}\exp(-\lambda)+\gamma C_{q}\exp(-\lambda)}{C_{q,y}\exp(\lambda_{q})+(H-C_{q,y})\exp(-\lambda_{q})}+\gamma+s\frac{C_{q,y}\exp(\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}$ (97)
    $+\sum_{q^{\prime}\in{\mathcal{Q}},q^{\prime}\neq q}\frac{C_{q^{\prime}}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda)}.$ (98)
    It follows that $\lim_{\lambda\to\infty}\xi_{\tau}=s+\gamma$.
  • Otherwise, $\xi_{j}=\xi_{A,j}+\xi_{F,j}=\xi_{A,j}=s\frac{C_{j}\exp(-\lambda)}{C_{q,y}\exp(\lambda)+(H-C_{q,y})\exp(-\lambda_{q})}$.
    It follows that $\lim_{\lambda\to\infty}\xi_{j}=0$.
Then, Lemma D.2 follows directly from
$\lim_{s\to\infty}\lim_{\lambda\to\infty}\frac{\exp(\xi_{y})}{\exp(\xi_{y})+\exp(\xi_{\tau})+\sum_{x=1,x\neq y}^{N}\exp(\xi_{x})}$ $=\lim_{s\to\infty}\frac{\exp(s)}{\exp(s)+\exp(s+\gamma)+N-1}$
$=\lim_{s\to\infty}\frac{1}{1+\exp(\gamma)+(N-1)\exp(-s)}$
$=\frac{1}{1+\exp(\gamma)}$
$=1-\alpha\quad\text{as }\gamma\to\ln(\frac{\alpha}{1-\alpha}),$
and
$\lim_{s\to\infty}\lim_{\lambda\to\infty}\frac{\exp(\xi_{\tau})}{\exp(\xi_{y})+\exp(\xi_{\tau})+\sum_{x=1,x\neq y}^{N}\exp(\xi_{x})}$ $=\lim_{s\to\infty}\frac{\exp(s+\gamma)}{\exp(s)+\exp(s+\gamma)+N-1}$
$=\lim_{s\to\infty}\frac{\exp(\gamma)}{1+\exp(\gamma)+(N-1)\exp(-s)}$
$=\frac{\exp(\gamma)}{1+\exp(\gamma)}$
$=\alpha\quad\text{as }\gamma\to\ln(\frac{\alpha}{1-\alpha}).$
D.2 Analysis of Normalized Gradient Descent on Population and Empirical Losses
To avoid notational overload, we write $C^{(m)}_{q,y}$ for $C^{(m)}_{q,y^{(m)}}$. We first consider the case where $\alpha$ is known, and then extend the analysis to the case of unknown $\alpha$.
@@ D.2 Analysis of Normalized Gradient Descent on Population and Empirical Losses
D.2.1 Known $\alpha$: Running Normalized Gradient Descent on the Population Loss $L(\lambda)$
@@
By Lemma E.1, the derivative of $f(\lambda)=-C\lambda+\ln(\frac{e^{C\lambda}}{1-\alpha}+N-1)$ is negative. Hence, $\frac{dL_{\mathrm{pop}}}{d\lambda}<0$. It follows that running normalized gradient descent on $L_{\mathrm{pop}}(\lambda)$ from $\lambda_{0}=0$ gives
@@
D.2.2 Unknown $\alpha$: Proof Of Theorem 5.4
@@
By Lemma E.1, we have $\frac{dL_{\mathrm{emp}}}{d\lambda}<0$. As a result, running normalized gradient descent on $L_{\mathrm{emp}}$ gives $\lambda_{t}=\eta t$. The population loss is
@@
Appendix E Technical Lemmas
Lemma E.1.
@@ Appendix E Technical Lemmas
Appendix F Further Details on Experiments
F.1 Additional Details on the Experimental Setup
@@
We also experimented with several different values of learning rates ranging from $0.1$ to $0.8$. Consistent with the theoretical findings, we find that the more reparamterized a model is, the less sensitive it is to changes in the learning rate. All of our results are reported for learning rates set at either $0.1,0.2$ or $0.8$.
@@
Computing Resources The experiments are implemented in PyTorch. All experiments are run on a single-CPU computer. The processor is 11th Gen Intel(R) Core i7-11700K with $32$ GB RAM. Training all models simultaneously takes about $30$ minutes from start to finish.
@@
Figure 2: Population and Unseen Output Test Losses of Reparam-Linear-W with $\eta=0.1$ (first row), $\eta=0.5$ (second row). Population losses converge with limited generalization to unseen output words.
@@
F.2 Attention Layer Learns to Predict Output Tokens while Feed-Forward Layer Learns to Predict Noise Token
@@
To measure the extent to which we can separate the learning functionality of the attention layer and the feed-forward layer, we train three models Origin-Linear, Reparam-Linear and Reparam-Linear-W, and record the logits of each layer on the output tokens, the noise tokens and the maximum values in the logits of the two layers on all three noisy tasks with $\alpha=0.2,0.5$ and $0.8$. We use $\eta=0.1$ in all experiments. The results are reported in Figures 3 to 5.
@@
We say that the attention layer and the feed-forward layer learns to predict the output and noise tokens, respectively, if $\xi_{A,y}=\max_{j}\xi_{A,j}$ and $\xi_{F,\tau}=\max_{j}\xi_{F,\tau}$. It can be observed that all three models exhibit some layer-specific learning mechanism, including the original model where all three matrices ${\bm{V}},{\bm{W}}$ and $\bm{F}$} are trained from scratch without reparameterization. However, depending on the noise level, at least one of the two layers in the original model do not fully specialize in either the output nor the noise tokens. For $\alpha\leq 0.5$, Figures 3(a) and 3(d) show that the attention layer in Origin-Linear learns to predict output tokens perfectly, however the feed-forward layer does not always predict $\tau$. At $\alpha=0.8$, the feed-forward layer succeeds in learning to predict $\tau$ but the attention layer fails to focus entirely on the output tokens.
@@
In contrast, the fully-reparameterized model Reparam-Linear exhibits perfect separation in the functionality of the two layers, which verified our Theorem 5.6. The same phenomenon is also observed in Reparam-Linear-W, which suggests that reparameterizing $\bm{F}$} is sufficient to force the two layers to be biased towards two different types of tokens.
@@
Figure 3: Origin-Linear with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.
@@
Figure 4: Reparam-Linear-W with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.
@@
Figure 5: Reparam-Linear with $\alpha=0.2$ (first row), $\alpha=0.5$ (second row) and $\alpha=0.8$ (third row). The learning rate is $\eta=0.1$.